OpenStudy (anonymous):

A 2.00g of sodium hydroxide, NaOH (the major ingredient in the drain cleaner) was dissolved in water to produce a total volume of exactly 200mL of solution. what is the molar concentration of this sodium hydroxide solution? (atomic mass of Na= 23.0g, o=16g and H = 1.0g)

OpenStudy (abhisar):

Hello @dinisha !

OpenStudy (abhisar):

Molar concentration or molarity = \[\frac{ Moles~of~solute }{ Volume~of~solution }\]

OpenStudy (abhisar):

So, here we can see 1 mole of NaOH = 23+16+1=40 g

OpenStudy (abhisar):

Now in 2 g there will be 1/20 moles

OpenStudy (abhisar):

Putting this and the given volume in the above formula we get that the molarity is 0.25 M

OpenStudy (anonymous):

thank u @Abhisar :)

OpenStudy (abhisar):

\(\Huge\text{Anytime !}\) \(\huge\ddot\smile\)